OpenStudy (ilovecake):

What is the answer? [(12 − 5) • 12] ÷ 6

OpenStudy (dumbcow):

12 - 5 = 7 12 /6 = 2 --> 7*2 = 14

OpenStudy (anonymous):

(7x12)÷6=?

OpenStudy (midhun.madhu1987):

Please follow this order: Bracket Division Multiplication Addition Subtraction

OpenStudy (ilovecake):

Ok thanks to both of you. :)

OpenStudy (anonymous):

14

OpenStudy (lieutenantgeneral):

www.wolframalpha.com